sveit) (c//usrta ds av (s au Ghne running-carf loop lulen verwed fan the east The fat S due lor ar a s Tne s The lagnihe
Answer #1

The magnitude of torque on the loop will be:

\tau = \mu Bsin\theta

=> \tau = N [\pi r^2 I] Bsin90\degree

N = 200

so, \tau = 200 [\pi (0.45)^2 95] (3.25\times 10^{-5})sin90\degree = 0.3928 Nm .